# Relativistic interacting fluids in cosmology

### v1

@@ I Introduction
In standard inflationary cosmologies, there are whole series of events in the very early universe (pre- and reheatings, baryo- and leptogeneses, etc.) which call for various particle number violations that can be described by well elaborated, and, at least in the case of nucleosynthesis, well established, particle physics theories Linde (1990); Davidson et al. (2008); Allahverdi et al. (2010); Marzola and Raidal (2016); Karam et al. (2020). In late-universe cosmology, and particularly in the context of the dark energy problem, a plethora of models have been proposed that introduce non-minimal interactions between dark energy and (dark or visible) matter Farrar and Peebles (2004); Pourtsidou et al. (2013); Wang et al. (2016); Carrillo González and Trodden (2018); Teixeira (2024), featuring conformal Amendola (2000); Koivisto (2005); Wetterich (2007); Kase and Tsujikawa (2020a); Poulot et al. (2024), disformal Zumalacarregui et al. (2013); van de Bruck and Morrice (2015); Cembranos and Maroto (2016); Teixeira et al. (2020); van de Bruck and Teixeira (2020), 3-form Ngampitipan and Wongjun (2011); Koivisto and Nunes (2013); Barreiro et al. (2017); Yao et al. (2018); Yao and Meng (2020), geometric Koivisto (2006); Harko and Lobo (2020); Lobo and Harko (2022); Boehmer and del Sordo (2024); Iosifidis and Pallikaris (2024), elastic Asghari et al. (2019); Figueruelo et al. (2021); Beltrán Jiménez et al. (2021a); Poulin et al. (2023); Cardona and Figueruelo (2022) or merely parameterised Valiviita et al. (2008); Clemson et al. (2012); Yang et al. (2018a, b); Ferlito et al. (2022) couplings. Currently, a strong motivation to consider such interactions is their potential to alleviate the existing tensions in the cosmological data, especially the Hubble tension Odderskov et al. (2016); Di Valentino et al. (2020); Gómez-Valent et al. (2020); Hoerning et al. (2023); Pan and Yang (2023).
@@
The aim of this article is to develop a universal and physically sound framework for the description of interacting fluids. Even in the case that the model of the dark sector is underpinned by an elementary particle theory [^1] An interesting current example is the scenario wherein a fraction of QCD axion dark matter is converted by a novel level crossing mechanism into dark energy, which is provided by a pseudo-Nambu Goldstone boson exploiting the temperature-scaling of its mass in a confining phase transition Müürsepp et al. (2024); Müürsepp (2024)., the collective description of the microscopic degrees of freedom at the level of hydrodynamics can be useful in order to understand the implications of the model to cosmological large scale structures. As will be shown in this article, an opposite proceeding is also possible, to wit, a consistent quantum field theory can emerge from a phenomenological model of thermodynamical processes by integrating out variables from the relativistic fluid action functional.
@@ II Hydrodynamical interactions
These definitions agree with [^4] In the notation and terminology of Misner, Thorne & Wheeler Misner et al. (1973) inherited by Brown Brown (1993), our $h$ is denoted $\mu$ and called the chemical potential and our $\mu$ is denoted $f$ and called the chemical free energy. We rather use the conventions of standard textbooks on thermodynamics. In Ref. Misner et al. (1973) the first law was deduced assuming (baryon) number conservation, and it may have often went unnoticed that actually (2) is the universal form of the law. We will show this explicitly later at (43).
@@
The (non-)conservation laws obtained by the variation of $I$ wrt the momenta $\varphi$ and $\theta$ are [^5] Here and in the following we have assumed that the interaction vectors $X^{\mu}$, $Y^{\mu}$ do not depend on the fluid variables, but the generalisation with $X(\varphi,\theta)$, $Y(\varphi,\theta)$ is straightforward. We will encounter (effective) generalisations in terms of kinetic terms for the momenta in Section IV and in terms of potential terms for the momenta in Section V. However, the interaction vectors should not depend upon $n$ or $s$ lest the interpretation of $\varphi$ and $\theta$ is modified.
@@
where $\textrm{D}_{\mu}$ is the torsion-free and metric-compatible Levi-Civita covariant derivative. The variations wrt the particle flux and the entropy per particle result, respectively, in the equations
@@
These relatioons are as foundational as (2) in our formulation and it is crucial that they are not violated despite the possible non-adiabatic or non-conservative properties of the system.
The variations wrt the Lagrange multipliers $\beta_{A}$ yield
@@
The variations wrt to the metric (and the connection, in the generalisation considered in Section IV) determine the gravitational sources. From the metric variations we obtain
@@
where the fluid energy-momentum tensor and the interaction energy-momentum tensor are given by [^7] The interactions are associated with energy-momentum, except in the (exotic) case that the densities $\mathcal{X}^{\mu}$ and $\mathcal{Y}^{\mu}$ are metric-independent (“levitons”).
@@
The projection of the metric-covariant divergence of the fluid energy-momentum tensor along the fluid flow gives
@@ III Scalar-fluid models
Here we consider the fluid interacting with a scalar field $\phi$. Scalar-fluid momentum exhange Pourtsidou et al. (2013); Boehmer et al. (2015a, b) is a consistent and interesting possibility in cosmology Koivisto et al. (2015); Dutta et al. (2018); Kase and Tsujikawa (2020b, c); Amendola and Tsujikawa (2020). Our new formulation is based on general thermodynamic principles, but the main interest in the new models is that they have different observational implications.
@@ III.1 Algebraic interactions
The non-conservation laws (7) follow directly as
@@ III.2 Derivative interactions
$\text{energy density}:\quad\hat{\rho}{}^{(\phi)}$ $=$ $\frac{1}{2}\left(\partial\phi\right)^{2}+V+X^{\prime}\phi_{,\alpha}\varphi^{,\alpha}+Y^{\prime}\phi_{,\alpha}\theta^{,\alpha}\,,$ (21a)
$\text{pressure}:\quad\hat{p}{}^{(\phi)}$ $=$ $\frac{1}{2}\left(\partial\phi\right)^{2}-V+X^{\prime}\phi_{,\alpha}\varphi^{,\alpha}+Y^{\prime}\phi_{,\alpha}\theta^{,\alpha}\,,$ (21b)
[…]
Unlike in the algebraic case, the energetics of the interactions can not be described simply in terms of an effective potential. The fluid equations now involve second derivatives of the scalar field,
@@ IV Hyperhydrodynamical models
$X_{\mu\nu}+Y_{\mu\nu}$ $=$ $-2\left[Q_{,(\mu)}\partial_{,\nu}+g_{\mu\nu}\left(\nabla^{\alpha}\nabla_{\alpha}+T^{\alpha}\partial_{\alpha}-\bar{Q}^{\alpha}\partial_{\alpha}\right)\right]f^{(1)}$ (29a)
@@
For a geometrical interpretation, the hypermomentum can be decomposed into
@@ IV.1 Palatini General Relativity
$I=\int\textrm{d}^{4}x\sqrt{-g}\left[\frac{m_{P}^{2}}{2}R-\rho(\lvert J\rvert,s)+\left(a_{1}Q^{\mu}+a_{2}\bar{Q}^{\mu}+\frac{8a_{1}+2a_{2}}{3}T^{\mu}\right)f_{,\mu}+\mathcal{J}^{\mu}\left(\varphi_{,\mu}+s\theta_{,\mu}+\beta_{A}\alpha^{A}{}_{,\mu}\right)\right]\,,$ (33)
@@ V Cosmology
Finally, we will consider the cosmological implications of the interacting models. We take gravity to be given by the usual General Relativity and adapt the fluid to a flat FLRW geometry, so that $u^{\mu}=\delta^{\mu}_{0}$, $h_{0\mu}=0$ and $h_{ij}=a^{2}(t)\delta_{ij}$, where the latin indices denote the spatial components. The (non-)conservation equations (7) then read
@@
and the energy-momentum non-conservation laws (13) reduce to one non-trivial non-continuity equation
@@ V.1 A phase transition of a gas of dark matter-energy particles (emerging from a 3-form)
and as long as $\phi_{0}/m\gg\mu t$, it is indistinguishable from a cosmological constant $\Lambda=m^{2}\phi_{0}^{2}/2$. The density of the interacting dust fluid evolves according to $\dot{\rho}+3H\rho=\mu m^{3}\left(\phi_{0}-m\mu t\right)$, which in the regime $\phi_{0}/m\gg\mu t$ mimics the CDM evolution with an effective source term. If we consider that the mass scale $m$ is of the order of the present Hubble rate and $\phi_{0}$ is of the order of the Planck mass, the source term will be negligible for a wide range of the parameter $\mu$. By adjusting the dark matter mass $\mu$, the effect of the coupling may become non-negligible between the last scattering and the present, providing the subtle boost in the dark matter density needed to address the Hubble tension.
Figure 1: Evolution of the cosmological equation of state $w$ as a function of $\log{a}$ for four different parameter choices is plotted in dashed lines, and the cosmological equation of state $\Lambda$CDM model which is plotted in solid line for comparison. Left panel: the model with $\mu>0$, with examples of initial conditions that give similar cosmology as the $\Lambda$CDM model but distinguishable from it. Right panel: the model with $\mu<0$. In one of the cases, the evolution of the cosmological equation of state coincides with the $\Lambda$CDM model. The three other examples feature $w<-1$ and are clearly distinguishable from the $\Lambda$CDM model.
The action for this model has the same number of parameters as the $\Lambda$CDM model, given now by the two mass parameters $\mu$ and $m$. The physical interpretation of the model is that of self-interacting matter. By integrating out the auxiliary scalar field, the action reads, explicitly in terms of the fundamental fields,
@@
Thus, we have single fluid of particles with mass $\mu$ which are undergoing chemical processes energy scale is $m$. In the limit $m\rightarrow 0$, the fluid becomes pressuless dust (CDM), and in the limit $\mu\rightarrow 0$ the fluid becomes vacuum energy ($\Lambda$). Since this dust-$\Lambda$ fluid isn’t associated with a temperature, we could drop the variables $s$ and $\theta$ without loss of generality in terms of the dynamical solutions. Let us do this, integrate out also the chemical momentum $\varphi$, perform a partial integration and rescale the vector $J^{\mu}\rightarrow J^{\mu}/m^{2}$. This procedure reveals that the thermodynamical fluid model is underpinned by a simple vector field theory,
@@
where $F_{\alpha\beta\gamma\delta}=4\textrm{D}_{[\alpha}A_{\beta\gamma\delta]}$ is the field strength of the 3-form. It is well-known that the massless 3-form $m^{2}\mu=0$ provides an effective $\Lambda$-term, and our result suggests that by adding a potential which is linear in the norm can, in some dynamical regime, provide an effective dust component besides the $\Lambda$-term. In the case of ideal dust the background 3-form components $A_{ijk}=a^{3}\epsilon_{ijk}m^{-2}n$ would be strictly constant, since they describe the comoving density of the dust. However, the dynamics is more subtle and the evolution of the components $A_{ijk}$ cannot be neglected even in the $\Lambda$CDM-like regime when the Hubble rate is small enough, as seen from the expression for the cosmological equation of state
@@
Finally, one could consider flipping the sign of the 3-form potential $V=\mu m^{2}\sqrt{A^{2}/24}$. Then the cosmological equation could become phantom-like $w<-1$, as seen from its expression equivalent to (49),
@@
In the fluid picture, the unconventional sign would seem to correspond to negative mass of the gas of particles, $\mu<0$, and the action would not remain unbounded from below as the number density of the particles $n\rightarrow\infty$. Most likely, this case would be pathological, as usual for models with phantom-like $w<-1$. However, validity of the naive stueckelbergisation argument that $V^{\prime}<0$ implies a ghost Koivisto and Nunes (2009) is not obvious for the special case of linear potential. For a canonical mass term $V=m_{C}^{2}A^{2}/2$ one easily establishes the equivalence of the 3-form with a scalar field that would be a ghost in the case that the 3-form mass term is tachyonic $m_{C}^{2}<0$ Koivisto and Nunes (2009). However, the equivalence with a scalar field breaks down for the linear potential $V=\mu m^{2}\sqrt{A^{2}/24}$. For curiosity, we plot some examples of the cosmological evolution of $w$ in the right hand side panel of Figure 1. It is curious that also this (more dubious) model can mimic the $\Lambda$CDM cosmology. Since the result $c_{S}^{2}=0$ holds regardless of the sign of the potential, at least classical instabilities are absent.
We recall that a dust 3-form appears in the Lorentz gauge theory as an energy source in the effective Einstein equations Złośnik et al. (2018). A massive space arises as a quantum effect in the sense that quantum fluctuations do not satisfy the Hamiltonian constraint of classical General Relativity and may contribute an apparent dust source Gallagher et al. (2024). The dust-like 3-form only exists in a symmetry-broken phase of the Lorentz gauge theory, whereas in the model (48) the square root forces an analogous symmetry breaking. It is then tempting to speculate that the model (48) could be an effective description of quantum gravity at cosmological scales, as it seems plausible for the kinetic term of the 3-form to be induced by quantum corrections. However, the possible embedding of the model into a fundamental theory of spacetime and gravity, as well as the detailed assessment of the model’s compatibility with the data, are outside the scope of the present article.

### v2

@@ I Introduction
In standard inflationary cosmologies, there are whole series of events in the very early universe (pre- and reheatings, baryo- and leptogeneses, etc.) which call for various particle number violations that can be described by well elaborated, and, at least in the case of nucleosynthesis, well established, particle physics theories Linde (1990); Davidson et al. (2008); Allahverdi et al. (2010); Marzola and Raidal (2016); Karam et al. (2020). In late-universe cosmology, and particularly in the context of the dark energy problem, a plethora of models have been proposed that introduce non-minimal interactions between dark energy and (dark or visible) matter Farrar and Peebles (2004); Pourtsidou et al. (2013); Wang et al. (2016); Carrillo González and Trodden (2018); Teixeira (2024), featuring conformal Amendola (2000); Koivisto (2005); Wetterich (2007); Kase and Tsujikawa (2020a); Poulot et al. (2024), disformal Zumalacarregui et al. (2013); van de Bruck and Morrice (2015); Cembranos and Maroto (2016); Teixeira et al. (2020); Vagnozzi et al. (2020); van de Bruck and Teixeira (2020), 3-form Ngampitipan and Wongjun (2011); Koivisto and Nunes (2013); Barreiro et al. (2017); Yao et al. (2018); Yao and Meng (2020), geometric Koivisto (2006); Harko and Lobo (2020); Lobo and Harko (2022); Boehmer and del Sordo (2024); Iosifidis and Pallikaris (2024), elastic Asghari et al. (2019); Figueruelo et al. (2021); Beltrán Jiménez et al. (2021a); Poulin et al. (2023); Cardona and Figueruelo (2022) or merely parameterised Valiviita et al. (2008); Clemson et al. (2012); Yang et al. (2018a, b); de Cesare and Wilson-Ewing (2022); Ferlito et al. (2022); Nunes et al. (2022) couplings. Currently, a strong motivation to consider such interactions is their potential to alleviate the existing tensions in the cosmological data, especially the Hubble tension Odderskov et al. (2016); Di Valentino et al. (2020a, b); Gómez-Valent et al. (2020); Hoerning et al. (2023); Pan and Yang (2023).
@@
The aim of this article is to develop a universal and physically sound framework for the description of interacting fluids. Phenomenological models without a proper Lagrangian formulation are compromised due to their arbitrariness or lack of predictivity, as well as unphysical features like instabilities they can introduce (e.g. Valiviita et al. (2008); Herrera et al. (2016); Yang et al. (2018a); Aljaf et al. (2020); de Cesare and Wilson-Ewing (2022)). Even in the case that the model of the dark sector is based on an elementary particle theory [^1] An interesting current example is the scenario wherein a fraction of QCD axion dark matter is converted by a novel level crossing mechanism into dark energy, which is provided by a pseudo-Nambu Goldstone boson exploiting the temperature-scaling of its mass in a confining phase transition Müürsepp et al. (2024); Müürsepp (2024)., the collective description of the microscopic degrees of freedom at the level of a hydrodynamical Lagrangian can be useful in order to understand the implications of the model to cosmological large scale structures. As will be shown in this article, an opposite proceeding is also possible, to wit, a consistent quantum field theory can emerge from a phenomenological model of thermodynamical processes by integrating out variables from the relativistic fluid action functional.
@@ II Hydrodynamical interactions
These definitions agree with [^4] In the notation and terminology of Misner, Thorne & Wheeler Misner et al. (1973) inherited by Brown Brown (1993), our $h$ is denoted $\mu$ and called the chemical potential and our $\mu$ is denoted $f$ and called the chemical free energy. We rather use the conventions of standard textbooks on thermodynamics. In Ref. Misner et al. (1973) the first law was deduced assuming (baryon) number conservation, and it may have often went unnoticed that actually (2) is the universal form of the law, as will be shown this explicitly later at (44). In terms of the entropy density $ns$ the first law would read $\textrm{d}\rho=T\textrm{d}(ns)+\mu\textrm{d}n$, consistently with e.g. Ballesteros et al. (2016); Luongo and Muccino (2018).
@@
The (non-)conservation laws obtained by the variation of $I$ w.r.t. the momenta $\varphi$ and $\theta$ are [^5] Here and in the following we have assumed that the interaction vectors $X^{\mu}$, $Y^{\mu}$ do not depend on the fluid variables, but the generalisation with $X(\varphi,\theta)$, $Y(\varphi,\theta)$ is straightforward. We will encounter (effective) generalisations in terms of kinetic terms for the momenta in Section IV and in terms of potential terms for the momenta in Section V. However, the interaction vectors should not depend upon $n$ or $s$ lest the interpretation of $\varphi$ and $\theta$ is modified.
@@
where $\textrm{D}_{\mu}$ is the torsion-free and metric-compatible Levi-Civita covariant derivative. The variations w.r.t. the particle flux and the entropy per particle result, respectively, in the equations
@@
These relations are as foundational as (2) in our formulation and it is crucial that they are not violated despite the possible non-adiabatic or non-conservative properties of the system.
The variations w.r.t.the Lagrange multipliers $\beta_{A}$ yield
@@
The variations w.r.t. to the metric (and the connection, in the generalisation considered in Section IV) determine the gravitational sources. From the metric variations we obtain
@@
where the fluid energy-momentum tensor and the interaction energy-momentum tensors are given by [^7] The interactions are associated with energy-momentum, except in the (exotic) case that the densities $\mathcal{X}^{\mu}$ and $\mathcal{Y}^{\mu}$ are metric-independent “levitons” (if some fields don’t couple to the metric at all, they don’t “gravitate” but “levitate”).
@@
The result (12a) confirms the consistency of the nomenclature in (1a,1e). The projection of the metric-covariant divergence of the fluid energy-momentum tensor along the fluid flow gives
@@ III Scalar-fluid models
Here we consider the fluid interacting with a scalar field $\phi$. Scalar-fluid momentum exhange Pourtsidou et al. (2013); Boehmer et al. (2015a, b) is a consistent and interesting possibility in cosmology Koivisto et al. (2015); Dutta et al. (2018); Kase and Tsujikawa (2020b, c); Amendola and Tsujikawa (2020). In our new formulation, it follows from the principles of thermodynamics in a generic situation wherein the phase space structure of the $\rho$-fluid may depend on the $\phi$-field ($Y^{\mu}\neq 0$) or the $\phi$-particles may decay into $\rho$-particles or vice versa ($X^{\mu}\neq 0$). The main interest in the new models is that they have different observational implications.
@@ III.1 Algebraic interactions
The (non-)conservation laws (7) follow directly as
@@ III.2 Derivative interactions
$\text{energy density}:\quad\hat{\rho}{}^{(\phi)}$ $=$ ${-}\frac{1}{2}\left(\partial\phi\right)^{2}+V+X^{\prime}\phi_{,\alpha}\varphi^{,\alpha}+Y^{\prime}\phi_{,\alpha}\theta^{,\alpha}\,,$ (21a)
$\text{pressure}:\quad\hat{p}{}^{(\phi)}$ $=$ ${-}\frac{1}{2}\left(\partial\phi\right)^{2}-V+X^{\prime}\phi_{,\alpha}\varphi^{,\alpha}+Y^{\prime}\phi_{,\alpha}\theta^{,\alpha}\,,$ (21b)
$\text{{momentum} flux vector}:\quad\hat{q}_{\mu}^{(\phi)}$ $=$ $\frac{1}{\sqrt{-(\partial\phi)^{2}}}\left(X^{\prime}\phi_{,\alpha}\varphi^{,\alpha}+Y^{\prime}\phi_{,\alpha}\theta^{,\alpha}\right)\phi_{,\mu}+\sqrt{-(\partial\phi)^{2}}\left(X^{\prime}\varphi_{,\mu}+Y^{\prime}\theta_{,\mu}\right)\,.$ (21c)
Unlike in the algebraic case, the energetics of the interactions can not be described simply in terms of an effective potential. If we assume that the fluid is comoving with the scalar field, the above expressions assume a somewhat simpler and more transparent form. Note that the assumption always holds in a Friedmann-Lemaître-Robertson-Walker background, and can be adopted as a legitimate gauge choice at the level of cosmological perturbations. We can then write
$\hat{\rho}{}^{(\phi)}$ $=$ $\rho^{(\phi)}+\sqrt{-(\partial\phi)^{2}}\left(X^{\prime}\mu+Y^{\prime}T\right)\,,$ (22a)
$\hat{p}{}^{(\phi)}$ $=$ $p^{(\phi)}+\sqrt{-(\partial\phi)^{2}}\left(X^{\prime}\mu+Y^{\prime}T\right)\,,$ (22b)
$\hat{q}_{\mu}^{(\phi)}$ $=$ $\sqrt{-(\partial\phi)^{2}}\left(X^{\prime}\varphi_{,\mu}+Y^{\prime}\theta_{,\mu}\right)+\left(X^{\prime}\mu+Y^{\prime}T\right)u_{\mu}\,.$ (22c)
The fluid equations now involve second derivatives of the scalar field,
@@ IV Hyperhydrodynamical models
$X_{\mu\nu}+Y_{\mu\nu}$ $=$ $-2\left[Q_{,(\mu}\partial_{,\nu)}+g_{\mu\nu}\left(\nabla^{\alpha}\nabla_{\alpha}+T^{\alpha}\partial_{\alpha}-\bar{Q}^{\alpha}\partial_{\alpha}\right)\right]f^{(1)}$ (30a)
@@
For a kinematical interpretation, the hypermomentum can be decomposed into
@@ IV.1 Palatini General Relativity
$I=\int\textrm{d}^{4}x\sqrt{-g}\left[\frac{m_{P}^{2}}{2}R-\rho(\lvert J\rvert,s)+\left(a_{1}Q^{\mu}+a_{2}\bar{Q}^{\mu}+\frac{8a_{1}+2a_{2}}{3}T^{\mu}\right)f_{,\mu}+{J}^{\mu}\left(\varphi_{,\mu}+s\theta_{,\mu}+\beta_{A}\alpha^{A}{}_{,\mu}\right)\right]\,,$ (34)
@@ V Cosmology
Finally, we will consider the cosmological implications of the interacting models. We take gravity to be given by the usual General Relativity and adapt the fluid to a flat Friedmann-Lemaître-Robertson-Walker geometry, so that $u^{\mu}=\delta^{\mu}_{0}$, $h_{0\mu}=0$ and $h_{ij}=a^{2}(t)\delta_{ij}$, where the latin indices denote the spatial components. The (non-)conservation equations (7) then read
@@
and the energy-momentum conservation laws (13) reduce to one non-trivial continuity equation
@@ V.1 A phase transition of a gas of dark matter-energy particles (emerging from a 3-form)
and as long as $\phi_{0}/m\gg\mu t$, it is indistinguishable from a cosmological constant $\Lambda=m^{2}\phi_{0}^{2}/2$. The density of the interacting dust fluid evolves according to $\dot{\rho}+3H\rho=\mu m^{3}\left(\phi_{0}-m\mu t\right)$, which in the regime $\phi_{0}/m\gg\mu t$ mimics the CDM evolution with an effective source term. If we consider that the mass scale $m$ is of the order of the present Hubble rate and $\phi_{0}$ is of the order of the Planck mass, the source term will be negligible for a wide range of the parameter $\mu$. By adjusting the dark matter mass $\mu$, the effect of the coupling may become non-negligible between the last scattering and the present, providing a boost which in the dark matter density which could potentially address the Hubble tension.
Figure 1: Evolution of the cosmological equation of state $w$ (left panel) and the fractional effective dark energy density $\Omega=\hat{V}/3m_{P}^{2}H^{3}$ (right panel) as a function of the e-folding time $\log{a}$ for four parameters and initial conditions chosen such that the four cases exhibit some qualitatively different (not necessarily realistic) features available in the simple model. The thick solid line reproduces the exactly the standard $\Lambda$CDM cosmology. The dashed line shows a faster transition to acceleration. The two other scenarios, plotted with thin solid and thin dotted lines, are characterised by a transition to a phantom-like cosmological equation of state $w<-1$, which is accompanied by effective $\Omega>1$ that implies a negative effective dark matter density, as explained in the text.
The action for this model has the same number of parameters as the $\Lambda$CDM model, the magnitude of $\Lambda$ and CDM now apparently traded for the two mass parameters $\mu$ and $m$. A physical interpretation of the model is that of self-interacting matter. By integrating out the auxiliary scalar field, the action reads, explicitly in terms of the fundamental fields,
@@
Thus, we have single fluid of particles with mass $\mu$ which are undergoing chemical processes energy scale is $m$. In the limit $m\rightarrow 0$, the fluid behaves as pressureless dust (CDM), and in the limit $\mu\rightarrow 0$ the fluid behaves as vacuum energy ($\Lambda$). Since this dust-$\Lambda$ fluid isn’t associated with a temperature, we could drop the variables $s$ and $\theta$ without loss of generality in terms of the dynamical solutions. Let us do this, integrate out also the chemical momentum $\varphi$, perform a partial integration and rescale the vector $J^{\mu}\rightarrow J^{\mu}/m^{2}$. This procedure reveals that the thermodynamical fluid model is underpinned by a simple vector field theory,
@@
where $F_{\alpha\beta\gamma\delta}=4\textrm{D}_{[\alpha}A_{\beta\gamma\delta]}$ is the field strength of the 3-form. It is well-known that the massless 3-form $m^{2}\mu=0$ provides a $\Lambda$-term Duff and van Nieuwenhuizen (1980); Hawking (1984); Weinberg (1989), and our result suggests that a potential which is linear in the norm the 3-form can, at least in some dynamical regime, provide also a CDM-term besides the $\Lambda$-term. In the case of ideal dust the background 3-form components $A_{ijk}=a^{3}\epsilon_{ijk}m^{-2}n$ would be strictly constant, since they describe the comoving density of the dust. However, the dynamics is more subtle and the evolution of the components $A_{ijk}$ cannot be neglected even in the $\Lambda$CDM-like regime when the Hubble rate is small enough, as seen from the expression for the cosmological equation of state
@@
Finally, we note that the sign of the 3-form potential can be flipped by the dynamics. Then the cosmological equation could become phantom-like $w<-1$, as seen from its expression equivalent to (50),
@@
The latter formula is in terms of the fraction $\Omega_{A}$ of the 3-form energy and the total energy of the universe. In the fluid picture, the unconventional sign would seem to correspond to negative energy density of the gas of particles, and the action would not remain unbounded from below as the number density of the particles $\lvert n\rvert\rightarrow\infty$. This case could be expected to be pathological, as usual for models with phantom-like $w<-1$. However, validity of the naive stueckelbergisation argument that $V^{\prime}<0$ implies a ghost Koivisto and Nunes (2009) is not obvious for the special case of linear potential. In addition, whilst for a canonical mass term $V=m_{C}^{2}A^{2}/2$ one easily establishes the equivalence of the 3-form with a massive scalar field that would be a ghost in the case that the 3-form mass term is tachyonic $m_{C}^{2}<0$ Koivisto and Nunes (2009) but, however, the equivalence with a scalar field breaks down for the linear potential $V=\mu m^{2}\sqrt{A^{2}/24}$. Thus, two standard arguments against $w<-1$ break down for the special case we have arrived at. For curiosity, we plot two examples of cosmological evolution of $w$ in Figure 1 also in cases such that $w<-1$ occurs. It is interesting that also this (more dubious) model can mimic the $\Lambda$CDM cosmology. Since the result $c_{S}^{2}=0$ holds regardless of the sign of the potential, at least classical instabilities are absent.
We recall that a dust 3-form appears in the Lorentz gauge theory as an energy source in the effective Einstein equations Złośnik et al. (2018). A massive space arises as a quantum effect in the sense that quantum fluctuations do not satisfy the Hamiltonian constraint of classical General Relativity and may contribute an apparent dust source Gallagher et al. (2024). The dust-like 3-form only exists in a symmetry-broken phase of the Lorentz gauge theory, whereas in the model (49) the square root forces an analogous symmetry breaking. It is then tempting to speculate that the model (49) could be an effective description of quantum gravity at cosmological scales, as it seems plausible for the kinetic term of the 3-form to be induced by quantum corrections. Indeed, the 3-form appears linearly in the extension of the Einstein-Cartan action, and there is no reason why the term could not have either sign. However, the possible embedding of the model into a fundamental theory of spacetime and gravity, as well as the detailed assessment of the model’s compatibility with the data, are outside the scope of the present article.